Nelson Corp is considering the purchase of a new piece of equipment. The cost savings from the equipment would result in an annual increase in cash flow of $173,850. The equipment will have an initial cost of $605,000 and have a 5 year life. If the salvage value of the equipment is estimated to be $250,000, what is the accounting rate of return? Ignore income taxes.

a. 28.74%

b. 17.00%

c. 14.50%

d. 30.41%
Answer #1

Depreciation expense per year = (605000-250000/5) = 71000

Net income = 173850-71000 = 102850

Accounting rate of return = Annual net income/initial investment = 102850/605000 = 17.00%

So answer is b) 17.00%